OpenStudy (candyapple123):

y=4x-1 and y=-1/4x+1000 is it parallel, perpendicular or neither

OpenStudy (anonymous):

You can answer that if you knowthe relationship between the slopes of two perpendicular lines.

OpenStudy (anonymous):

perpendicular :4 multiply by -1/4 is -1

OpenStudy (anonymous):

\[m_1: slope line1\] \[m_2: slope line2\]

OpenStudy (anonymous):

if \[m_1\times m_2 = -1\] then the lines are perpendicular.

OpenStudy (anonymous):

\[m_1 = 4\] and \[m_2 = -\frac{1}{4}\]

OpenStudy (anonymous):

So, \[m_1\cdot m_2 = -4\cdot\frac{1}{4}=-1\]

OpenStudy (anonymous):

Perpendicular

OpenStudy (anonymous):

Did you get that?

OpenStudy (candyapple123):

yeah

OpenStudy (anonymous):

medal then?